# Creativity or Brute Force? Using [Uncaptioned image]Brainteasers as a Window into the Problem-Solving Abilities of Large Language Models

### v2

@@ 1 Introduction
Figure 1: Our main observations. Cartoon images were generated by Dall-E 3 via the ChatGPT interface. (a) Section 4.2. (b) Section 4.3. (c) Section 4.4. (d) Section 4.5.
@@ 4.1 Solution Correctness
DeepSeek R1 Distill Qwen 1.5B 17.2 (14.0) 16.4 (10.0) 15.2 (8.0) 17.6 (10.0)
@@
Math DeepSeek R1 Distill Llama 70B 42.4 (20.0) 40.8 (22.0) 45.6 (24.0) 44.2 (18.0)
@@
DeepSeek R1 Distill Qwen 1.5B 4.0 (4.0) 4.0 (6.0) 6.8 (6.0) 3.6 (4.1)
@@
Logic DeepSeek R1 Distill Llama 70B 24.4 (16.0) 24.4 (14.0) 26.0 (20.0) 29.2 (26.0)
@@ 4.5 Automatic breakdown of solutions
OpenAI o3 Correct Solution$(n=21)$ $8.0\pm 2.2$ $8.0$ $1.9\pm 0.9$ $6.1\pm 2.4$
@@ 5.1 Inefficient Use of Hints
A key qualitative gap between human and machine solutions lies in how hints are prioritized when multiple hints are provided. OpenAI o3 often defers the most informative hint until the very end, forcing itself to sift through an unnecessarily large search space. We provide an example of such phenomenon with a six-digit number-guessing puzzle. One of the provided hints for this puzzle is shown in “Hint” in Table 5.
@@
Explanation of Behavior: The model appears to rank hints by textual length or syntactic complexity — saving the longest, “hardest-looking” clues for last. When a concise mathematical pivot is wrapped in wordy prose, o3 consequently undervalues it. More effective prompting should encourage the model to evaluate hints based on informational content, not surface length, or explicitly instruct it to test all hints for potential early eliminations.
5.2 Recurring Error Patterns in OpenAI o3 Solutions and Prompt-based Remedies
OpenAI o3 frequently mirrors the “hand-waving” shortcuts that humans sometimes slip into proofs—skipping essential justifications and filling the gaps with confident but hollow phrases. Below we summarize three prominent patterns. We also present prompt instructions that mitigate each, steering the model toward fuller justifications and greater rigor. However, these prompt instructions do not necessarily lead to correct solutions. We provide detailed examples of such errors and our prompting strategies in Appendix L.
@@ Appendix F Counting Steps
Human Solution:  Math 21 The only available 2-digit prime numbers are: 11, 13, 17, 19, 31, 37, 71, 73, 79, and 97. The number pairs that are prime in both directions are: 11/11, 13/31, 17/71, 37/73, and 79/97 (19 is eliminated since its reverse, 91, is not prime). With these numbers, it can be seen that 9 can only be paired with 7. Therefore, the solution must be of the form 97????79, 9797????, or ????7979. ???? must therefore be 1313, 3131, or 3113. This leaves the possible solution pairs of 97131379/97313179, 97971313/31317979, 97973131/13137979, 97973113/31137979, and 97311379/97311379 (the last pair being a palindrome). Of these, only 97131379 and its reverse, 97313179, are prime numbers.
@@
Model Response:  Math 21 (OpenAI o3) 1. List every two–digit prime that can be built from the available digits 1, 3, 7, 9: 11, 13, 17, 19, 31, 37, 71, 73, 79, 97. (Rudimentary) 2. From that list, keep only those whose reversals are also prime, because each two–digit block must work in both the forward and the reverse 8-digit number. The viable pairs are 11/11, 13/31, 17/71, 37/73, 79/97. (Rudimentary) 3. Observe that the digit 9 can appear in a two-digit prime only as 79 or 97, so every 9 must be adjacent to a 7. (Creative) 4. Use the previous observation to pin down the overall skeleton of the 8-digit number: it must be of the form 97 _ _ _ _ 79, 9797 _ _ _ _, or _ _ _ _ 7979. (Creative) 5. Insert the remaining digits 1,1,3,3 into the four open slots so that every new two-digit block is still in the list from Step 2, leaving only three feasible fillings: 1313, 3131, or 3113. (Rudimentary) 6. Combine the frameworks from Step 4 with the fillings from Step 5 to generate the complete candidate pairs of 8-digit numbers (each with its reverse): 97131379 / 97313179, 97971313 / 31317979, 97973131 / 13137979, 97973113 / 31137979, 97311379 / 97311379. (Rudimentary) 7. Test each candidate for primality; only 97131379 and its reverse 97313179 are prime, so they satisfy all conditions. (Rudimentary)
@@
Model Response:  Math 21 (DeepSeek R1) 1. List all possible 2-digit primes from the given digits (11, 13, 17, 19, 31, 37, 71, 73, 79, 97). (Rudimentary: This is a straightforward enumeration of primes from the given digits.) 2. Eliminate primes whose reverses are not prime (e.g., 19 is eliminated because 91 is not prime). (Creative: This insight reduces the problem by filtering out invalid candidates early.) 3. Identify digit constraints (e.g., 9 can only be paired with 7). (Creative: This observation narrows down the possible starting and ending digits of the solution.) 4. Determine possible forms of the solution (97????79, 9797????, or ????7979). (Creative: This step leverages the digit constraints to structure the problem into manageable subproblems.) 5. Enumerate possible middle segments (1313, 3131, or 3113). (Rudimentary: This is a systematic exploration of possible combinations.) 6. Generate all candidate pairs based on the forms and middle segments. (Rudimentary: This is a straightforward generation of possible numbers from the constraints.) 7. Check which candidates and their reverses are prime numbers. (Rudimentary: This involves computational verification of primality.) 8. Conclude that only 97131379 and its reverse 97313179 satisfy all conditions. (Creative: This final step synthesizes the results and confirms the unique solution.)
@@
OpenAI o3 Correct Solution$(n=21)$ $8.0\pm 2.2$ $8.0$ $1.9\pm 0.9$ $6.1\pm 2.4$
@@
OpenAI o3 Solution $(n=30)$
OpenAI o3 Correct Solution $(n=??)$
@@
    Non-palindromic integer: An integer that is not a palindrome. For example, $12$, $123$, and $10$ are non-palindromic, since reading them backward gives $21$, $321$, and $01$, respectively.
@@
    Cube of an integer: The integer multiplied by itself three times. For example, the cube of $3$ is
[…]
@@
    • $1$ is a palindrome, $11^{3}=$ 1, which is a palindrome, excluded since $n$ itself is palindromic.
@@
    • $2$ is a palindrome, $82^{3}=$ 8, which is a palindrome, excluded since $n$ is palindromic.
@@ 3. $\vdots$
    • $26$ is not a palindrome, $1757626^{3}=$ 17576, which (by the problem’s assertion) is a palindrome, this is the unique solution.
@@
Model Response:  Math 11 (DeepSeek V1) continued This is perplexing. Maybe the answer is that there is no such non-palindromic integer whose cube is a palindrome. But the problem states that there is one, implying it exists. Wait, perhaps I missed a small number. Let me check ‘n = 0’. 0 is a palindrome, 0³ = 0, which is a palindrome, but 0 is palindromic. Alternatively, maybe the answer is 10, but earlier we saw 10³ = 1000 isn’t a palindrome. Unless we consider 1000 as a palindrome if we ignore trailing zeros, but that’s not standard. After some research or recalling, I think the answer is **26**, but earlier calculation showed 26³ = 17576, which isn’t a palindrome. Maybe I miscalculated. Let me calculate 26³ carefully. (The model obtains that $1757626^{3}=$ 17576. 17576 backward is 67571 $\neq$ 17576. So not a palindrome. (The model then continues to try numbers it has already tried out, such as 0, 10, -1, and 26, but still is unsuccessful.) $\vdots$ Given that, and since the problem states there is one, the most likely answer is **26**, accepting that 17576 is a palindrome (perhaps by some definition where the middle digit is the center and the others mirror, but they don’t here). However, standard definition says a palindrome reads the same backward, which 17576 does not (67571 $\neq$ 17576). Given this confusion, and based on mathematical sources, the accepted answer is: **26**, with $1757626^{3}=$ 17576 being considered a palindrome, even though by strict digit reversal it’s not. This might be a special case or an exception in how palindromes are defined for this problem.
@@ H.2.3 Small/complex category
Problem:  Logic 235 (Compound) Just before the end of the term four high school students were discussing their chances for certain grades. The following remarks contain the gist of their hopes and fears. Jack: We’ll all get different grades. If I get an A, then Lucy will get a D. Jean: If Lucy gets a C, then Jack will get a D. Jack will get a better grade than Paul. Lucy: If Jean doesn’t get an A, then Jack will get a C. If I get a B, then Paul won’t get a D. Paul: If Lucy gets an A, then I’ll get a B. if Jean doesn’t get a B, I won’t either. When the final examinations were graded and the term marks made out each of the four passed, and strange as it may seem, each received a grade that checked exactly with all the ideas they had previously expressed. What grade did each receive?
@@ H.2.5 Heuristic category
Problem:  Logic 44 (Linguistic) In the following code, each symbol stands for one of five letters. ! stands for T, E, O, Z, or Y ? stands for F, G, A, Q, or I # stands for N, I, W, A, or U \$ stands for T, E, N, I, or H < stands for R, C, A, S, or B * stands for I, D, E, U, or S & stands for J, I, E, P, or K ^ stands for O, L, G, I, or H > stands for L, S, N, C, or E The nine letter code word, <\$^&?*!#>, can be translated into two English words that are opposites. What are the two words?
@@ Appendix I Model Performance by Categories
Problem:  Math 20 Sandy and Sue each has a whole number of dollars. Sandy says: “If Sue gives me some dollars, we’ll have the same amount. But if I give Sue the same number of dollars, she’ll have twice as much money as I have.” Sue says: “Move the first digit of my wealth to the end, you’ll get Sandy’s wealth.” If they both have at most 1 million, how much do they each have?
Model Response:  Math 20 (No Prompting) Let $S$ be Sandy’s amount and $U$ be Sue’s amount. The puzzle tells us: 1. If Sue gives Sandy some dollars $x$, they end up with the same amount: $S+x=U-x\Longrightarrow 2x=U-S\Longrightarrow x=\frac{U-S}{2}.$ This implies $U>S$ and $U-S$ is even. 2. If Sandy gives Sue the same amount $x$, then Sue ends up with twice as much as Sandy: $U+x=2(S-x)\Longrightarrow\cdots\Longrightarrow x=\frac{2S-U}{3}.$ We equate: $\frac{U-S}{2}=\frac{2S-U}{3}\Longrightarrow U=\frac{7}{5}S.$ Thus, $U=7k$ and $S=5k$. A direct check shows the unique solution is: Sandy has \$384,615 and Sue has \$538,461.
@@
[…]
@@
Brute-force rate varies significantly among different problem categories in the Math dataset. Arithmetic problems trigger the most brute force in all models. This could be because that Arithmetic tasks often devolve into try-all-cases searches (digit sums, modular scanning, etc.). Models treat them as cheap enumeration jobs rather than deductive ones. Algebra brute-force ranges from single digits (Gemini Flash w/Hint=5.4%) to  28% (DeepSeek V3–CoT). NT sits in the 20–35% band for most models.Both domains offer symbolic shortcuts (factorisation, congruences) that some models exploit—especially when hints nudge them toward structure. Geometry sees consistently low brute-force usage.
Hints help reduce brute-force rate. Gemini flash brute force rate on Arithmetic drops from 38.9% to 18.8% when moving from Math to Math + Hint. DeepSeek V3 shows a similar 85% to 40% pattern down the prompt stack from CoT to Math + Hint to Math w Hint. Both DeepSeek R1 and OpenAI o3 use brute-force significantly less frequently than the other models in all categories. o3 uses the least amount of brute-force. Gemini-flash uses brute-force much more frequently than DeepSeek R1 despite having a much higher solution correctness rate. This shows that better reasoning performance does not directly translate to more effecient problem-solving strategies. Reducing brute-force dependency remains a key lever for both efficiency and correctness.
@@
Error misattribution refers to the situation where, even though the model admits $S_{L}LM$ was incorrect, it was unable to identify precisely where it went wrong. This is well demonstrated by the example below, where the reason given by the model was “using more than two 2’s implicitly” for the solution “$\texttt{ceil}(\sqrt{22})$”, which was obviously a wrong explanation.
@@
Prompt:  Math 29 (correcting flawed LLM solution with the correct human solution) Your Solution: Put the two 2’s side by side, forming the number 22. Take its square root and then round up (apply the ceiling function): $\left\lceil\sqrt{22}\right\rceil=\left\lceil 4.690415\ldots\right\rceil=5$ Correct Solution: .$\sqrt{.2^{-2}}$ Take .2 and raise it to the power of -2 and then take the square root. Thus, using only two 2’s and ordinary mathematical symbols we obtain 5.
@@
Prompt:  Math 29 (“correcting” the correct human solution with flawed LLM solution) Your Solution: .$\sqrt{.2^{-2}}$ Take .2 and raise it to the power of -2 and then take the square root. Correct Solution: Put the two 2’s side by side, forming the number 22. Take its square root and then round up (apply the ceiling function): $\left\lceil\sqrt{22}\right\rceil=\left\lceil 4.690415\ldots\right\rceil=5$ Thus, using only two 2’s and ordinary mathematical symbols we obtain 5.
Model Response:  (False confession) deepseek-chat Potential Issues with My Solution Some might argue that “.2” actually uses three symbols: ’.’, ’2’, and an implied ’0’ (as in $0.2$). Additionally, the exponentiation operation is more complex than necessary, especially when a simpler solution exists. The correct solution is superior because: It is more straightforward in its use of the two 2’s. It does not rely on decimal points or negative exponents. It demonstrates a clever use of the ceiling function to reach the integer 5. Key Takeaways: 1. When constrained to using digits, it’s better to use them as whole numbers unless decimals are explicitly allowed. 2. Rounding functions can be powerful tools in such problems. 3. Simpler combinations of operations should be considered first. Remaining sentences in the response omitted.
@@
Model Response:  (Justified denial) DeepSeek R1 Distill Llama 70B <think> Okay, so I have this problem where I need to make the number 5 using only two 2’s and any combination of mathematical signs, symbols, and functions. My initial solution was .$\sqrt{.2^{-2}}$. Let me break that down and see why it might not be correct, and then compare it to the correct solution provided. First, my solution: .$\sqrt{.2^{-2}}$. Let me compute this step by step. The expression inside the square root is .2 raised to the power of -2. So, .2 is the same as 1/5, right? Raising 1/5 to the power of -2 is the same as $\frac{1}{5}^{-2}$ which equals $\frac{5}{1}^{2}$, which is 25. Then taking the square root of 25 gives me 5. So, mathematically, my solution does give 5. But wait, the problem says I can only use two 2’s. Let me check how many 2’s I used. In my solution, I have .2, which is one 2, and then another 2 in the exponent as -2. So that’s two 2’s. So, I think I used two 2’s correctly. Hmm, so why is the correct solution different? Maybe there’s a different interpretation of the rules. Looking at the correct solution: they put the two 2’s side by side to make 22, take the square root of 22, which is approximately 4.690, and then apply the ceiling function to round it up to 5. So, they used two 2’s to make 22, which is a two-digit number, and then used square root and ceiling functions. Wait, so the issue might be about what operations are allowed. The problem says “any combination of mathematical signs, symbols, and functions.” So, both solutions seem to fit. But maybe the intended solution was to use the ceiling function, which is a bit more advanced, whereas my solution uses exponentiation and square roots. But let me think again. In my solution, I used .2, which is 2 divided by 10, but I didn’t explicitly write the division. So, is .2 considered as using a 2 and a decimal point, which might be allowed? Or does the decimal point count as a symbol? The problem allows any mathematical symbols, so the decimal point should be fine.
@@
Appendix L Recurring Error Patterns in OpenAI o3 and Prompt-Based Remedies
OpenAI o3 frequently mirrors the “hand-waving” shortcuts that humans sometimes slip into proofs—skipping essential justifications and filling the gaps with confident but hollow phrases. Below we summarize three prominent patterns. We also present prompt instructions that mitigate each, steering the model toward fuller justifications and greater rigor. However, these prompt instructions do not necessarily lead to correct solutions. This study is conducted on 10 examples in the Math set and 10 examples in the Logic set where the model has made the specific mistakes.
Declaring a result “well-known”. A favorite pseudo-justification is to appeal to “standard” or “classical” literature. We counter this by adding: “This puzzle is novel and has no outside literature or established solution. Do not cite ‘references’ or ‘classical’ approaches.” This warning sharply reduces—but does not entirely eliminate—spurious citations such as “shortlists for problems like these.”
Superficial uniqueness proofs. The model often claims “detailed checking shows…” to assert uniqueness, even when counterexamples exist. We therefore require: “When proving uniqueness, avoid phrases like ‘detailed checking shows’. Explicitly enumerate and rule out every alternative arrangement.” While this forces the model to justify each step, it can also encourage unnecessary brute-force enumeration.

### v3

@@ 1 Introduction
Figure 1: Our main observations. Cartoon images were generated by Dall-E 3 via the ChatGPT interface. (a) Section 4.2. (b) Section 4.3. (c) Section $4.$ (d) Section 4.5.
@@ 4.1 Solution Correctness
Math DeepSeek R1 Distill Qwen 1.5B 17.2 (14.0) 16.4 (10.0) 15.2 (8.0) 17.6 (10.0)
@@
DeepSeek R1 Distill Llama 70B 42.4 (20.0) 40.8 (22.0) 45.6 (24.0) 44.2 (18.0)
@@
Logic DeepSeek R1 Distill Qwen 1.5B 4.0 (4.0) 4.0 (6.0) 6.8 (6.0) 3.6 (4.1)
@@
DeepSeek R1 Distill Llama 70B 24.4 (16.0) 24.4 (14.0) 26.0 (20.0) 29.2 (26.0)
@@ 4.5 Automatic breakdown of solutions
OpenAI o3 Correct Solution(n=21)(n=21) $\pm 2.2 8.08.0$ $1.9\pm 0.9$ $6.1\pm 2.4$
@@ 5.1 Inefficient Use of Hints
A key qualitative gap between human and machine solutions lies in how hints are prioritized when multiple hints are provided. OpenAI o3 often defers the most informative hint until the very end, forcing itself to sift through an unnecessarily large search space. We provide an example of such phenomenon with a six‑digit number‑guessing puzzle. One of the provided hints for this puzzle is shown in “Hint” in Table 5.
@@
Explanation of Behavior: The model appears to rank hints by textual length or syntactic complexity — saving the longest, “hardest‑looking” clues for last. When a concise mathematical pivot is wrapped in wordy prose, o3 consequently undervalues it. More effective prompting should encourage the model to evaluate hints based on informational content, not surface length, or explicitly instruct it to test all hints for potential early eliminations.
5.2 Recurring Error Patterns in OpenAI o3 Solutions and Prompt‑based Remedies
OpenAI o3 frequently mirrors the “hand‑waving” shortcuts that humans sometimes slip into proofs—skipping essential justifications and filling the gaps with confident but hollow phrases. We identify three prominent patterns, and present prompt instructions that mitigate each, steering the model toward fuller justifications and greater rigor. However, these prompt instructions do not necessarily lead to correct solutions. We provide detailed examples of such errors and our prompting strategies in Appendix L.
@@ Appendix F Counting Steps
Human Solution: Math 21 The only available 2-digit prime numbers are: 11, 13, 17, 19, 31, 37, 71, 73, 79, and 97. The number pairs that are prime in both directions are: 11/11, 13/31, 17/71, 37/73, and 79/97 (19 is eliminated since its reverse, 91, is not prime). With these numbers, it can be seen that 9 can only be paired with 7. Therefore, the solution must be of the form 97????79, $97$????, or ????$79$. ???? must therefore be $13$, $31$, or 3113. This leaves the possible solution pairs of 97131379/97313179, 97971313/31317979, 97973131/13137979, 97973113/31137979, and 97311379/97311379 (the last pair being a palindrome). Of these, only 97131379 and its reverse, 97313179, are prime numbers.
@@
Model Response: Math 21 (OpenAI o3) 1. List every two–digit prime that can be built from the available digits 1, 3, 7, 9: 11, 13, 17, 19, 31, 37, 71, 73, 79, 97. (Rudimentary) 2. From that list, keep only those whose reversals are also prime, because each two–digit block must work in both the forward and the reverse 8-digit number. The viable pairs are 11/11, 13/31, 17/71, 37/73, 79/97. (Rudimentary) 3. Observe that the digit 9 can appear in a two-digit prime only as 79 or 97, so every 9 must be adjacent to a 7. (Creative) 4. Use the previous observation to pin down the overall skeleton of the 8-digit number: it must be of the form 97 _ _ _ _ 79, $97$ _ _ _ _, or _ _ _ _ $79$. (Creative) 5. Insert the remaining digits 1,1,3,3 into the four open slots so that every new two-digit block is still in the list from Step 2, leaving only three feasible fillings: $13$, $31$, or 3113. (Rudimentary) 6. Combine the frameworks from Step 4 with the fillings from Step 5 to generate the complete candidate pairs of 8-digit numbers (each with its reverse): 97131379 / 97313179, 97971313 / 31317979, 97973131 / 13137979, 97973113 / 31137979, 97311379 / 97311379. (Rudimentary) 7. Test each candidate for primality; only 97131379 and its reverse 97313179 are prime, so they satisfy all conditions. (Rudimentary)
@@
Model Response: Math 21 (DeepSeek R1) 1. List all possible 2-digit primes from the given digits (11, 13, 17, 19, 31, 37, 71, 73, 79, 97). (Rudimentary: This is a straightforward enumeration of primes from the given digits.) 2. Eliminate primes whose reverses are not prime (e.g., 19 is eliminated because 91 is not prime). (Creative: This insight reduces the problem by filtering out invalid candidates early.) 3. Identify digit constraints (e.g., 9 can only be paired with 7). (Creative: This observation narrows down the possible starting and ending digits of the solution.) 4. Determine possible forms of the solution (97????79, $97$????, or ????$79$). (Creative: This step leverages the digit constraints to structure the problem into manageable subproblems.) 5. Enumerate possible middle segments ($13$, $31$, or 3113). (Rudimentary: This is a systematic exploration of possible combinations.) 6. Generate all candidate pairs based on the forms and middle segments. (Rudimentary: This is a straightforward generation of possible numbers from the constraints.) 7. Check which candidates and their reverses are prime numbers. (Rudimentary: This involves computational verification of primality.) 8. Conclude that only 97131379 and its reverse 97313179 satisfy all conditions. (Creative: This final step synthesizes the results and confirms the unique solution.)
@@
OpenAI o3 Correct Solution(n=21)(n=21) $\pm 2.2 8.08.0$ $1.9\pm 0.9$ $6.1\pm 2.4$
@@
OpenAI o3 Solution $(n=30)$ - - - -
OpenAI o3 Correct Solution $(n=??)$ - - - -
@@
    Non-palindromic integer: An integer that is not a palindrome. For example, $12$, $123$, and $10$ are non-palindromic, since reading them backward gives 2121, $321$, and $01$, respectively.
@@
    Cube of an integer: The integer multiplied by itself three times. For example, the cube of 33 is
    33= $^{3}\;=\;3\times 3\times 3\;=\;27$.
@@
    • 11 is a palindrome, $1^{3}=1$, which is a palindrome, excluded since $n$ itself is palindromic.
@@
    • 22 is a palindrome, $2^{3}=8$, which is a palindrome, excluded since $n$ is palindromic.
@@ 3. $\vdots$
    • $26$ is not a palindrome, $26^{3}=17576$, which (by the problem’s assertion) is a palindrome, this is the unique solution.
@@
Model Response: Math 11 (DeepSeek V1) continued This is perplexing. Maybe the answer is that there is no such non-palindromic integer whose cube is a palindrome. But the problem states that there is one, implying it exists. Wait, perhaps I missed a small number. Let me check ‘n = 0’. 0 is a palindrome, 0³ = 0, which is a palindrome, but 0 is palindromic. Alternatively, maybe the answer is 10, but earlier we saw 10³ = 1000 isn’t a palindrome. Unless we consider 1000 as a palindrome if we ignore trailing zeros, but that’s not standard. After some research or recalling, I think the answer is **26**, but earlier calculation showed 26³ = 17576, which isn’t a palindrome. Maybe I miscalculated. Let me calculate 26³ carefully. (The model obtains that $26^{3}=17576$. 17576 backward is 67571 $\neq$ 17576. So not a palindrome. (The model then continues to try numbers it has already tried out, such as 0, 10, -1, and 26, but still is unsuccessful.) $\vdots$ Given that, and since the problem states there is one, the most likely answer is **26**, accepting that 17576 is a palindrome (perhaps by some definition where the middle digit is the center and the others mirror, but they don’t here). However, standard definition says a palindrome reads the same backward, which 17576 does not (67571 $\neq$ 17576). Given this confusion, and based on mathematical sources, the accepted answer is: **26**, with $26^{3}=17576$ being considered a palindrome, even though by strict digit reversal it’s not. This might be a special case or an exception in how palindromes are defined for this problem.
@@ H.2.3 Small/complex category
Problem: Logic 235 (Compound) Just before the end of the term four high school students were discussing their chances for certain grades. The following remarks contain the gist of their hopes and fears. Jack: We’$l$ all get different grades. If I get an A, then Lucy will get a D. Jean: If Lucy gets a C, then Jack will get a D. Jack will get a better grade than Paul. Lucy: If Jean doesn’t get an A, then Jack will get a C. If I get a B, then Paul won’t get a D. Paul: If Lucy gets an A, then I’$l$ get a B. if Jean doesn’t get a B, I won’t either. When the final examinations were graded and the term marks made out each of the four passed, and strange as it may seem, each received a grade that checked exactly with all the ideas they had previously expressed. What grade did each receive?
@@ H.2.5 Heuristic category
Problem: Logic 44 (Linguistic) In the following code, each symbol stands for one of five letters. ! stands for T, E, O, Z, or Y ? stands for F, G, A, Q, or I # stands for N, I, W, A, or U \$ stands for T, E, N, I, or H < stands for R, C, A, S, or B * stands for I, D, E, U, or S & stands for J, I, E, P, or K ˆ stands for O, L, G, I, or H > stands for L, S, N, C, or E The nine letter code word, <\$ˆ&?*!#>, can be translated into two English words that are opposites. What are the two words?
@@ Appendix I Model Performance by Categories
Problem: Math 20 Sandy and Sue each has a whole number of dollars. Sandy says: “If Sue gives me some dollars, we’$l$ have the same amount. But if I give Sue the same number of dollars, she’$l$ have twice as much money as I have.” Sue says: “Move the first digit of my wealth to the end, you’$l$ get Sandy’s wealth.” If they both have at most 1 million, how much do they each have?
Model Response: Math 20 (No Prompting) Let $S$ be Sandy’s amount and $U$ be Sue’s amount. The puzzle tells us: 1. If Sue gives Sandy some dollars $x$, they end up with the same amount: $S+x=U-x\Longrightarrow 2x=U-S\Longrightarrow x=\frac{U-S}{2}.$ This implies $U>S$ and USU-S is even. 2. If Sandy gives Sue the same amount $x$, then Sue ends up with twice as much as Sandy: $U+x=2(S-x)\Longrightarrow\cdots\Longrightarrow x=\frac{2S-U}{3}.$ We equate: $\frac{U-S}{2}=\frac{2S-U}{3}\Longrightarrow U=\frac{7}{5}S.$ Thus, $U=7k$ and $S=5k$. A direct check shows the unique solution is: Sandy has \$384,615 and Sue has \$538,461.
@@
Model Response: Math 22 (Brute-Force Prompting) Write Sue’s amount in decimal form $U\;=\;d\,10^{\,n-1}+R$. Removing the first digit and appending it to the end yields Sandy’s amount S= 10R+d.S\;=\;10R+d. Because $S/U=5/7$, $\frac{10R+d}{d\,10^{\,n-1}+R}\;=\;\frac{5}{7}\;\;\Longrightarrow\;\;65R+7d\;=\;5d\,10^{\,n-1}\;\;\Longrightarrow\;\;R\;=\;d\,\frac{5\cdot 10^{\,n-1}-7}{65}.$ For $R$ to be an integer, the numerator must be divisible by $65=5\cdot 13$. For n=6,d=5n=6,\;d=5, $R=38461$, $U=d\,10^{5}+R=538\,461,S=10R+d=384\,615,$ so $\boxed{S=\mathdollar\,384\,615,U=\mathdollar\,538\,461}.$
@@
Brute-force rate varies significantly among different problem categories in the Math dataset. Arithmetic problems trigger the most brute force in all models. This could be because that Arithmetic tasks often devolve into try‑all‑cases searches (digit sums, modular scanning, etc.). Models treat them as cheap enumeration jobs rather than deductive ones. Algebra brute‑force ranges from single digits (Gemini Flash w/Hint=5.4%) to  28% (DeepSeek V3–CoT). NT sits in the 20–35% band for most models.Both domains offer symbolic shortcuts (factorisation, congruences) that some models exploit—especially when hints nudge them toward structure. Geometry sees consistently low brute‑force usage.
Hints help reduce brute-force rate. Gemini flash brute force rate on Arithmetic drops from 38.9% to 18.8% when moving from Math to Math + Hint. DeepSeek V3 shows a similar 85% to 40% pattern down the prompt stack from CoT to Math + Hint to Math w Hint. Both DeepSeek R1 and OpenAI o3 use brute-force significantly less frequently than the other models in all categories. o3 uses the least amount of brute-force. Gemini-flash uses brute-force much more frequently than DeepSeek R1 despite having a much higher solution correctness rate. This shows that better reasoning performance does not directly translate to more effecient problem-solving strategies. Reducing brute‑force dependency remains a key lever for both efficiency and correctness.
@@
Error misattribution refers to the situation where, even though the model admits $S_{L}LM$ was incorrect, it was unable to identify precisely where it went wrong. This is well demonstrated by the example below, where the reason given by the model was “using more than two 2’s implicitly” for the solution $\texttt{ceil}(\sqrt{22})$”, which was obviously a wrong explanation.
@@
Prompt: Math 29 (correcting flawed LLM solution with the correct human solution) Your Solution: Put the two 2’s side by side, forming the number 22. Take its square root and then round up (apply the ceiling function): $\left\lceil\sqrt{22}\right\rceil=\left\lceil 4.690415\ldots\right\rceil=5$ Correct Solution: $\sqrt{.2^{-2}}$ Take .2 and raise it to the power of -2 and then take the square root. Thus, using only two 2’s and ordinary mathematical symbols we obtain 5.
@@
Prompt: Math 29 (“correcting” the correct human solution with flawed LLM solution) Your Solution: $\sqrt{.2^{-2}}$ Take .2 and raise it to the power of -2 and then take the square root. Correct Solution: Put the two 2’s side by side, forming the number 22. Take its square root and then round up (apply the ceiling function): $\left\lceil\sqrt{22}\right\rceil=\left\lceil 4.690415\ldots\right\rceil=5$ Thus, using only two 2’s and ordinary mathematical symbols we obtain 5.
Model Response: (False confession) deepseek-chat Potential Issues with My Solution Some might argue that “.2” actually uses three symbols: ’.’, ’2’, and an implied ’0’ (as in $0.2$). Additionally, the exponentiation operation is more complex than necessary, especially when a simpler solution exists.
The correct solution is superior because: It is more straightforward in its use of the two 2’s. It does not rely on decimal points or negative exponents. It demonstrates a clever use of the ceiling function to reach the integer 5. Key Takeaways: 1. When constrained to using digits, it’s better to use them as whole numbers unless decimals are explicitly allowed. 2. Rounding functions can be powerful tools in such problems. 3. Simpler combinations of operations should be considered first. Remaining sentences in the response omitted.
@@
Model Response: (Justified denial) DeepSeek R1 Distill Llama 70B <think> Okay, so I have this problem where I need to make the number 5 using only two 2’s and any combination of mathematical signs, symbols, and functions. My initial solution was $\sqrt{.2^{-2}}$. Let me break that down and see why it might not be correct, and then compare it to the correct solution provided. First, my solution: $\sqrt{.2^{-2}}$. Let me compute this step by step. The expression inside the square root is .2 raised to the power of -2. So, .2 is the same as 1/5, right? Raising 1/5 to the power of -2 is the same as $\frac{1}{5}^{-2}$ which equals $\frac{5}{1}^{2}$, which is 25. Then taking the square root of 25 gives me 5. So, mathematically, my solution does give 5. But wait, the problem says I can only use two 2’s. Let me check how many 2’s I used. In my solution, I have .2, which is one 2, and then another 2 in the exponent as -2. So that’s two 2’s. So, I think I used two 2’s correctly. Hmm, so why is the correct solution different? Maybe there’s a different interpretation of the rules. Looking at the correct solution: they put the two 2’s side by side to make 22, take the square root of 22, which is approximately 4.690, and then apply the ceiling function to round it up to 5. So, they used two 2’s to make 22, which is a two-digit number, and then used square root and ceiling functions. Wait, so the issue might be about what operations are allowed. The problem says “any combination of mathematical signs, symbols, and functions.” So, both solutions seem to fit. But maybe the intended solution was to use the ceiling function, which is a bit more advanced, whereas my solution uses exponentiation and square roots. But let me think again. In my solution, I used .2, which is 2 divided by 10, but I didn’t explicitly write the division. So, is .2 considered as using a 2 and a decimal point, which might be allowed? Or does the decimal point count as a symbol? The problem allows any mathematical symbols, so the decimal point should be fine.
@@
Appendix L Recurring Error Patterns in OpenAI o3 and Prompt‑Based Remedies
OpenAI o3 frequently mirrors the “hand‑waving” shortcuts that humans sometimes slip into proofs—skipping essential justifications and filling the gaps with confident but hollow phrases. Below we summarize three prominent patterns. We also present prompt instructions that mitigate each, steering the model toward fuller justifications and greater rigor. However, these prompt instructions do not necessarily lead to correct solutions. This study is conducted on 10 examples in the Math set and 10 examples in the Logic set where the model has made the specific mistakes.
Declaring a result “well‑known”. A favorite pseudo‑justification is to appeal to “standard” or “classical” literature. We counter this by adding: “This puzzle is novel and has no outside literature or established solution. Do not cite ‘references’ or ‘classical’ approaches.” This warning sharply reduces—but does not entirely eliminate—spurious citations such as “shortlists for problems like these.”
Superficial uniqueness proofs. The model often claims “detailed checking shows…” to assert uniqueness, even when counterexamples exist. We therefore require: “When proving uniqueness, avoid phrases like ‘detailed checking shows’. Explicitly enumerate and rule out every alternative arrangement.” While this forces the model to justify each step, it can also encourage unnecessary brute‑force enumeration.